OpenStudy (anonymous):

What is an equation in point slope form of the line that passes through points (4,5) and (-3,-1)?

OpenStudy (anonymous):

Given: (4,5) = (x1, y1) and (-3,-1) = (x2, y2) So slope \[\huge m = \frac{y_1-y_2}{x_1-x_2} = \frac{5-(-1)}{4-(-3)} = \frac{5+1}{4+3} \] \[\huge \rightarrow m= \frac{6}{7} \] Equation in point slope form is given as: \[\huge y-y_1 = m(x-x_1)\] \[\huge y-5 = \frac{6}{7}(x-4)\] \[\huge 7( y-5) = 6(x-4)\] \[\huge 7 y-35 = 6x-24\] is the required eq in slope point form.

OpenStudy (anonymous):

@xxmaryyxx